Shift report 20011020 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Oct 20 2001 - 00:31:38 EDT

  • Next message: Apache: "Shift report 20011020 00:00-08:00"

    ------------------------------------------------------------
    Date: Sat Oct 20 00:31:37 2001, Shift: 16:00-24:00
    Supervisor: J.I.Jordre, with: 
    Log book#10 pages 36 - 37
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam from 17:48 to 18:39, when beam was lost because of blue
       quenc link interlock. Then problems with cryo Turbo expansion
       pump, trying to swith to backup pump, and finally reporting that
       there had been some oil leak in the pump so that the estimate is
       there will be no beam for at least 12 hours. I.e. probably no beam
       untill Saturday at 12 noon. If someone needs access to the IR, call
       MCR and they will go to restricted access.
    2. Run 5446	12k trigger 6, ended when I found BBL 37 had tripped
       Run 5447	32k trigger 6.
       TPC pedestal, DVM and pedestal runs 5448, 5449 and 5450 respectively.
    3. Only the mentioned trip of BBL 37.
       Turned down magnet currents and put them into stand-by mode before
       closing business.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sat Oct 20 2001 - 00:32:04 EDT